A ladder 5m is resting against vertical waall.The bottom of the ladder is pulled along the ground away from the wall at the rate of 1.5m/sec.The length of the highest point of the ladder when the foot of the ladder is 4.0m away from the wall decreases at the rate of

Grade:12

2 Answers

Pratham Ashish
14 years ago
hi, let the length of the highest point of the ladder = h distance of the bottom from wall = d so, h^2 + d^2 = 5^2 h^2 + d^2 = 25 h = (25 - d^2 )^1/2 dh/dt = - 1/2 (25 - d^2 )^ -1/2 * 2d = - 4 / 3 , rate of decreas of h 4/3 m/s
anandpal shekhawat
14 years ago
Hi, let a be vertical height & b be distance of bottom from wall, so a^2 + b^2 =25 when b = 4m then a = 3m & Diff. above w.r.t. dt, we can find speed of highest point